OpenStudy (anonymous):

find the factorization of the polynomial below. Enter each factor as a polynomial in descending order. 3x^2 + 7x + 2 This one confuses me.

OpenStudy (anonymous):

(3x+1)(x+2)

OpenStudy (anonymous):

Why is that though? I thought you had to make it equal to 7? or does it not matter?

OpenStudy (anonymous):

3x*2+x*1=6x+x=7x

OpenStudy (anonymous):

jahtoday is quite correct.. and you can check by using the FOIL method to multiply it out (3x+1)(x+2)=3x^2+7x+2

OpenStudy (anonymous):

Yes kenzie, using his numbers, they do add up to 7x

myininaya (myininaya):

if you have ax^2+bx+c, then all you have to do is find the factors of a*c that have product a*c and that add up to be b. a=3, b=7, c=2 a*c=3*2=6*1 b=7=6+1 so replace the bx=7x with 6x+1x 3x^2+7x+2 3x^2+6x+1x+2 and then factor by grouping 3x(x+2)+1(x+2) (x+2)(3x+1)

OpenStudy (anonymous):

It's tricky because this one adds a co-efficient to x^2 you'll have to pay mind to

OpenStudy (anonymous):

Which of the following polynomials does not have real roots? A.12x2 - 17x - 7 B.x4 + 5x2 + 6 C.2x4 + 13x3 + 21x2 D.x2 + x - 2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

How many roots does y = x2 - 6x + 9 have? It may help to graph the equation

OpenStudy (anonymous):

x^2